# Improving Long-Range Navigation with Spatially-Enhanced Recurrent Memory via End-to-End Reinforcement Learning

### v1

@@ 5 Experiments
To evaluate the effectiveness of the proposed Spatially-Enhanced Recurrent Unit (SRU) and the attention-based network architecture in enhancing long-horizon robot navigation, we conduct experiments in both simulated and real-world environments. We compare the SRU against standard LSTM and GRU units in long-range navigation tasks, focusing on their performance in end-to-end reinforcement learning (RL) training and navigation success rates (SR). Additionally, we compare the SRU policy, integrated with our proposed network structure and trained using recurrent RL, against the current state-of-the-art baseline (Lee et al., 2024) for legged-wheel robot navigation. The baseline employs an external mapping pipeline (Miki et al., 2022b) to integrate historical observations into a 2.5D grid map and uses an explicit historical path to address POMDP challenges in local navigation within unknown environments. Our evaluation highlights the advantages of the implicit recurrent memory provided by SRU in solving long-range navigation tasks across diverse environments.
Furthermore, we ablate the role of spatial attention layers in compressing features from encoded observations to improve memorization and overall navigation performance, contrasting this with the convolution and average pooling approach used in Wijmans et al. (2019). We also investigate the impact of regularization techniques on training the SRU unit end-to-end in RL, evaluating their effectiveness in preventing early convergence to suboptimal solutions and enhancing navigation performance. Finally, we explain and validate the pretrained image encoder’s ability to bridge the sim-to-real gap by demonstrating zero-shot transfer across diverse and complex real-world environments.
@@ 5.1 Experimental Setup
Method Maze Pillar Stair Pit Overall
@@
Table 1: Navigation success rate (SR) across four environment types: Maze, Random Pillars, Stairs, and Pits. The table compares standard LSTM and GRU units with our proposed spatially enhanced counterparts, including SRU-GRU, SRU-LSTM, and SRU-Ours.
@@ 5.2 Comparsion with Recurrent Units
We evaluate the performance of the proposed Spatially-Enhanced Recurrent Units (SRUs) compared to standard LSTM and GRU units. Given the superior spatial memorization capability of SRUs, as demonstrated in Figure 1, we hypothesize that integrating SRUs will improve the performance of navigation policies in addressing long-range navigation tasks. To test this hypothesis, we train networks integrated with different recurrent units end-to-end using RL in the simulated 3D environments shown in Figure 6 and evaluate their navigation performance.
@@
5.3 Comparsion against RL-based Navigation Baseline
@@
Next, we compare our proposed network structure, trained using recurrent reinforcement learning with SRU, against the state-of-the-art baseline method (Lee et al., 2024). The baseline employs an external mapping pipeline to integrate historical observations for local mapping (Miki et al., 2022b) and uses an explicit historical traversed path to address POMDP challenges in long-range navigation. While explicit mapping can theoretically achieve high accuracy in spatial-temporal registration, it has two major drawbacks: (i) it introduces significant delays that hinder real-time performance, especially on high-speed, agile platforms (Lee et al., 2024), and (ii) it relies on heuristic rules (e.g., fixed context window lengths) to select information, limiting its ability to capture complex spatial-temporal dependencies and abstract information beyond the selected context window. For instance, as shown in Figure 9(a), the baseline policy collects historical paths for approximately 20 meters, which is insufficient to navigate the long corridor spanning around 30 meters. In contrast, recurrent neural networks offer an unlimited context window and can learn intricate spatial-temporal dependencies optimized for the given task. This allows the SRU-based policy to adapt to long-horizon navigation challenges more effectively. Moreover, our end-to-end architecture processes raw depth sensor inputs, reducing latency and better supporting the agile and fast motion of legged-wheel platforms during deployment. To ensure consistency with the platform utilized in Lee et al. (2024), this comparison is conducted using the simulated wheeled ANYmal (Hutter et al., 2016) robot model, which differs from the robot model used in the RNN comparison experiment shown in Table 1. Both policies were trained under identical conditions and evaluated on an independent test environment set to ensure a fair comparison. Note that our policy relies solely on a front-facing camera with a limited field of view, whereas the baseline incorporates a local height scan with a similar range for environmental detection but benefits from a complete 360-degree field of view for mapping.
@@
The experimental results (see Table 2) indicate that our architecture, leveraging an implicit memory representation within the recurrent module, outperforms the baseline, achieving an overall 29.6% improvement in navigation SR. Figure 9(a) illustrates the representative comparison mentioned above: in a long-horizon maze environment, the baseline approach, constrained by the limited horizon of its explicit memory, fails to navigate through the maze and eventually loops in a long corridor. In contrast, our policy successfully traverses a dead-end corridor and reaches the goal, demonstrating the effectiveness of the SRU unit in learning implicit spatial-temporal mapping for long-horizon navigation tasks.
@@
Quantitatively, we analyze the success rate (SR) as a function of travel distance to evaluate long-range memorization and generalization capabilities. As shown in Figure 10, with a maximum episodic time of 60 seconds (consistent with training) and the robot’s maximum speed set to 1.5 m/s, the baseline approach’s SR drops significantly when the travel distance exceeds 40 meters. In contrast, our SRU-based approach maintains an SR of over 80% up to 50 meters. When the maximum episode time is extended to 120 seconds, the baseline’s SR still declines to below 60% at the same 40-meter distance, constrained by its fixed context window. Conversely, our SRU-based approach sustains an SR of over 70% up to 120 meters, demonstrating the SRU’s superior ability to implicitly memorize spatial-temporal mappings. This enables effective generalization to longer travel distances beyond the training range. The baseline’s reliance on a fixed explicit memory window limits its capacity to capture long-range dependencies, hindering its generalization in extended long-distance navigation tasks.
@@
Furthermore, we observed that the baseline policy struggles to effectively learn to climb staircases unless dense reward guidance is provided. This limitation arises from the difficulty explicit memory mechanisms face in capturing intricate spatial-temporal features, which are essential for the robot to develop the maneuvers required to overcome 3D obstacles effectively. Lastly, the end-to-end recurrent training setup offers a simpler and more maintainable solution compared to the baseline method. The baseline relies on an external mapping pipeline and the storage of additional historical paths for each robot, which can introduce complexity and overhead during both training and real-world deployment.
(a) Policy with Explicit Mapping and Historical Path
@@
Figure 9: Example comparison of navigation performance in a maze environment. The robot’s traversed trajectory is shown in yellow, with the traversal order indicated by A, B, C, and D. (a) The baseline approach starts looping in the long corridor between points B and C, failing to navigate through the dead-end corridors. (b) Our approach, with SRU recurrent memory, successfully navigates from start to goal, rerouting through the dead-end corridors and reaching the goal through area D.
Figure 10: Success rate sorted by travel distance: comparison between the baseline approach, which uses explicit mapping and a fixed-length historical path, and our approach, which employs the implicit recurrent memory of SRU. Our method maintains a high success rate over longer distances and extends effectively with longer episodic times. In contrast, the baseline’s success rate drops significantly for longer travel distances, even when the maximum episodic time is doubled, due to its fixed context window limitation.
@@
Model SR (%)
Baseline (Explicit Mapping/Path) 60.4
SRU-Ours (Implicit Recurrent Memory) 78.3
Table 2: Comparison of the overall navigation success rate (SR): The SRU-based implicit memory approach significantly outperforms the baseline RL method, which relies on explicit mapping and historical paths, in long-range navigation tasks.
Figure 12: Average return rewards during training: The network with the spatial attention layers achieves higher return compared to the network without the attention layers.
@@ 5.4 Importance of Spatial Attention Layers
We now examine the role of spatial attention layers in the network architecture and their impact on navigation performance. These layers are designed to compress and emphasize relevant features from encoded observations, addressing a key challenge faced by recurrent structures: the difficulty of retaining long-term information due to the exponential decay of memory over time. By selectively focusing on the most salient features, the attention mechanism emphasizes the most relevant spatial cues for navigation based on the robot’s state and reduces the information density passed into the recurrent memory at each step. We hypothesize that this mechanism can improve the network’s memorization and navigation capabilities, enabling it to handle complex, long-range tasks more effectively.
@@
To test this, we conduct an ablation study where we remove the attention layer from our network architecture and use convolution followed by average pooling to compress the encoded image features, as utilized in Wijmans et al. (2019). Figure 12 shows the average return rewards during training. The network without the attention layers exhibits significantly lower performance compared to the policy using the attention layers. Similarly, Table 3 shows the SR performance of policies trained with (w.) and without (w/o.) the attention layers. The results demonstrate that spatial attention layers significantly enhances the network’s navigation performance, achieving over a 56.2% improvement in terms of SR. Notably, the attention effect emerges naturally during the end-to-end RL training without requiring additional supervision or auxiliary losses. Figure 11 illustrates the attention weights generated by the cross-attention layer over raw visual inputs in three distinct real-world deployment scenarios: an indoor office, an outdoor terrace, and a forest environment. The attention weights, with four attention heads (depicted in different colors), dynamically emphasize the most relevant spatial cues, such as obstacles and navigable free space, based on the robot’s state at the time the depth input was recorded. This highlights the effectiveness of training the spatial attention mechanism end-to-end and its ability to generalize across diverse and challenging environments.
Model SR (%)
SRU-Ours w/o. Attention 50.5
SRU-Ours w. Attention 78.9
Table 3: Comparison of the overall navigation success rate (SR): The network with attention layers achieves a significantly higher SR compared to the network without attention layers, which uses only convolution followed by average pooling for spatial feature compression, as implemented in Wijmans et al. (2019).
@@ 5.6 Large-Scale Pretraining for Sim-to-Real Transfer
Table 4: Comparison of the overall navigation success rate (SR) with and without DML regularization for policies using LSTM and SRU units. DML significantly enhances SR for SRU (over 20%) and provides a marginal improvement for LSTM (2.8%), showcasing DML’s effectiveness in unlocking SRU’s potential for long-range navigation.
@@
Figure 18: Evaluation of long-range navigation in diverse real-world environments: (a) Main Hall of a university, (b) outdoor terrace, and (c) forest environment with natural obstacles. In each scenario, the robot is tasked with two separate navigation goals (memory reset between goals), resulting in two trajectory segments (orange and blue). Labels A, B, and C mark key areas traversed by the robot.
@@ 6 Limitations and Future Work
While the proposed SRUs in this paper demonstrate significant improvements in spatial-temporal learning capabilities, their recurrent nature remains subject to exponential memory decay, which can limit their ability to retain global context over extended sequences. As a result, the long-range navigation capabilities presented in this paper are primarily focused on local navigation using ego-centric sensing. Extending this approach to global navigation tasks—spanning kilometers or hours—may require additional mechanisms or architectural enhancements, such as integrating or maintaining a global map. Furthermore, while SRUs enhance the network’s capacity for implicit spatial memorization and improve long-range navigation performance, the precise characteristics of the information retained and utilized during end-to-end navigation training remain unclear. This highlights a broader challenge in explainable artificial intelligence, where understanding the internal representations and decision-making processes of neural networks continues to be an active area of research (Mi et al., 2024). Future work could explore integrating SRUs with advanced architectures, such as Transformers, to combine their strengths in handling long-range dependencies with the efficiency of recurrent structures. Investigating auxiliary losses or additional regularization techniques to further leverage the potential of spatial-temporal memorization in SRUs could also be beneficial. Additionally, extending the application of SRUs to other domains, such as robotic manipulation and 3D reconstruction, could unlock new possibilities and advancements in spatial-temporal learning. In summary, this paper highlights the potential of implicit spatial memory mechanisms in addressing complex navigation challenges while identifying opportunities for further exploration and optimization in both methodology and application.

### v2

@@ 5 Experiments
To evaluate the effectiveness of the proposed Spatially-Enhanced Recurrent Unit (SRU) and the attention-based network architecture in enhancing long-horizon robot navigation, we conduct experiments in both simulated and real-world environments. We compare the SRU against standard LSTM and GRU units in long-range mapless navigation tasks, focusing on their performance in end-to-end reinforcement learning (RL) training and navigation success rates (SR). Additionally, we compare the SRU policy, integrated with our proposed network structure and trained using recurrent RL, against two current state-of-the-art (SOTA) baselines (Huang et al., 2023; Lee et al., 2024) for robot navigation with RL. Our evaluation highlights the advantages of the implicit recurrent memory provided by SRU in solving long-range mapless navigation tasks across diverse environments.
Furthermore, we ablate the role of our proposed spatial attention layers in compressing features from encoded observations to improve memorization and overall navigation performance. We compare our approach against the convolution and average pooling method used in Wijmans et al. (2019), as well as the attention mechanism introduced in Huang et al. (2023). We also investigate the impact of regularization techniques on training the SRU unit end-to-end in RL, evaluating their effectiveness in preventing early convergence to suboptimal solutions and enhancing navigation performance. Finally, we explain and validate the pretrained image encoder’s ability to bridge the sim-to-real gap by demonstrating zero-shot transfer across diverse and complex real-world environments.
@@ 5.1 Experimental Setup
Model Maze Pillar Stair Pit Overall
@@
Table 1: Navigation success rate (SR) for policies integrated with different recurrent units across four environment types: Maze, Random Pillars, Stairs, and Pits. The table compares standard LSTM and GRU units with our proposed spatially enhanced counterparts: SRU-GRU, SRU-LSTM, and SRU-Ours.
@@ 5.2 Comparsion with Recurrent Units
We evaluate the performance of the proposed Spatially-Enhanced Recurrent Units (SRUs) compared to standard LSTM and GRU units. Given the superior spatial memorization capability of SRUs, as demonstrated in Figure 1, we hypothesize that integrating SRUs will improve the performance of navigation policies in addressing long-range mapless navigation tasks. To test this hypothesis, we train, under same conditions, policy networks integrated with different recurrent units end-to-end using RL in the simulated 3D environments shown in Figure 6. All policies are equipped with the same components (attention, training regularization, and a pretrained encoder); only the recurrent network structure differs. We then evaluate their navigation performance.
@@
5.3 Comparsion against RL-based Navigation Baselines
@@
Next, we compare our proposed network structure, trained using recurrent reinforcement learning with SRU, against two state-of-the-art RL baseline methods: the Goal-guided Transformer-based RL approach (GTRL) (Huang et al., 2023) and the RL approach with Explicit Mapping and Historical Path (EMHP) (Lee et al., 2024). GTRL employs a goal-guided Transformer (GoT) architecture to extract task-relevant visual features from stacked historical observations, enabling mapless navigation using only egocentric input. EMHP employs an external mapping pipeline to integrate historical observations for local mapping (Miki et al., 2022b) and uses an explicit historical traversed path to address POMDP challenges in long-range mapless navigation. While explicit mapping (EMHP) can theoretically achieve high accuracy in spatial-temporal registration, it has two major drawbacks: (i) it introduces significant delays that hinder real-time performance, especially on high-speed, agile platforms (Lee et al., 2024), and (ii) it relies on heuristic rules (e.g., fixed context window lengths) to select information, limiting its ability to capture complex spatial-temporal dependencies and abstract information beyond the selected context window.
Figure 9(a) presents a comparison between the EMHP baseline and our SRU-based approach. The EMHP policy collects historical paths for approximately 20 meters, which is insufficient to navigate the long corridor spanning around 30 meters. In contrast, recurrent neural networks offer an unlimited context window and can learn intricate spatial-temporal dependencies optimized for the given task. This allows the SRU-based policy to adapt to long-horizon navigation challenges more effectively. Moreover, our end-to-end architecture processes raw depth sensor inputs, reducing latency and better supporting the agile and fast motion of legged-wheel platforms during deployment. For the GTRL baseline, temporal history is provided by stacking several past observation frames, which are then fused using the transformer-based architecture as described in (Huang et al., 2023). Following the original approach in Huang et al. (2023), we use the 4-frame history in our experiments. However, the choice of the number of stacked frames remains heuristic: a short history may miss important context, while a longer history increases computational cost quadratically. For a fair comparison, we retrain the GTRL baseline, as described in Huang et al. (2023), within our environment. We replace the RGB input with depth images and utilize the same on-policy optimization method (PPO) for end-to-end RL training. To ensure consistency with the platform utilized in Lee et al. (2024), this comparison is conducted using the simulated wheeled ANYmal (Hutter et al., 2016) robot model, which differs from the robot model used in the other comparisons in this paper. All policies are trained under identical conditions and evaluated on an independent test environment set to ensure a fair comparison. Note that our policy and the GTRL baseline rely solely on a front-facing camera with a limited field of view, whereas the EMHP approach incorporates a local height scan with a similar range for environmental detection but benefits from a complete 360-degree field of view for mapping.
@@
The experimental results (see Table 2) indicate that our architecture, leveraging an implicit memory representation within the recurrent module, outperforms both baselines. Compared to EMHP, it achieves a 29.6% relative improvement in SR. Against GTRL, it demonstrates a remarkable 105.0% relative improvement. These results underscore the limitations of explicit mapping or stacked-frame histories for long-horizon navigation under limited context. Figure 9(a) illustrates a representative comparison: in a long-horizon maze environment, the EMHP approach, despite achieving a relatively higher SR among the two baselines, is constrained by the limited horizon of its explicit memory, fails to navigate through the maze and eventually loops in a long corridor. In contrast, our policy successfully traverses a dead-end corridor and reaches the goal, demonstrating the effectiveness of the SRU unit in learning implicit spatial-temporal mapping for long-horizon navigation tasks. To further validate the effectiveness of SRU for implicit spatial-temporal memorization in mapless navigation, we replaced GTRL’s stacked historical observations with our SRU-based recurrent memory while keeping all other components unchanged. This modified variant, denoted GTRL*, achieves a 73.6% relative improvement over the original GTRL baseline, increasing the SR from 38.2% to 66.3% (Table 2). This substantial gain highlights the advantage of SRU’s implicit recurrent memory over heuristic stacking of historical observations in capturing spatial-temporal dependencies for improved navigation performance. Notably, compared to GTRL*, our complete SRU-based approach (Ours) achieves an additional 18.1% relative improvement, attributable to our proposed spatial attention layers. The impact of these attention layers is further discussed later.
@@
Quantitatively, while the EMHP approach achieves comparable navigation performance, we also analyze the success rate (SR) as a function of travel distance to evaluate its long-range memorization and generalization capabilities. As shown in Figure 10, with a maximum episodic time of 60 seconds (consistent with training) and the robot’s maximum speed set to 1.5 m/s, the EMHP approach’s SR drops significantly when the travel distance exceeds 40 meters. In contrast, our SRU-based approach maintains an SR of over 80% up to 50 meters. When the maximum episode time is extended to 120 seconds, the EMHP’s SR still declines to below 60% at the same 40-meter distance, constrained by its fixed context window. Conversely, our SRU-based approach sustains an SR of over 70% for distances up to 120 meters, demonstrating the SRU’s superior ability to implicitly learn spatial-temporal mappings and generalize to distances beyond the training range. The baseline’s reliance on a fixed explicit memory window limits its capacity to capture long-range dependencies, hindering its generalization in extended long-distance navigation tasks.
@@
Furthermore, we observed that the EMHP policy struggles to effectively learn to climb staircases unless dense reward guidance is provided. We believe this limitation arises from the inherent difficulty explicit memory mechanisms face in capturing intricate spatial-temporal features, which are essential for the robot to develop the maneuvers required to overcome 3D obstacles effectively. Lastly, the end-to-end recurrent setup offers a simpler and more maintainable solution compared to baseline methods. In contrast, baselines rely on an external mapping pipeline and the storage of additional historical paths or observations for each robot, which can introduce complexity and overhead during both training and real-world deployment.
(a) Policy with Explicit Mapping and Historical Path (EMHP)
@@
Figure 9: Comparison of proposed mapless method with SRU recurrent memory against the EMHP baseline approach in a maze environment. The robot’s traversed trajectory is shown in yellow, with traversal order marked as A, B, C, and D. (a) The EMHP approach starts looping in the long corridor between points B and C, failing to navigate through the dead-end corridors. (b) Our approach, with SRU recurrent memory, successfully navigates from start to goal, rerouting through the dead-end corridors and reaching the goal through area D.
Figure 10: Success rate sorted by travel distance: comparison between the EMHP baseline approach, which uses explicit mapping and a fixed-length historical path, and our approach, which employs the implicit recurrent memory of SRU. Our method maintains a high success rate over longer distances and extends effectively with longer episodic times. In contrast, the baseline’s success rate drops significantly for longer travel distances, even when the maximum episodic time is doubled, due to its fixed context window limitation.
@@
Method SR (%)
GTRL (w. historical obs.) 38.2
EMHP (w. explicit mapping / path) 60.4
GTRL* (w. SRU memory) 66.3
Ours (w. SRU memory) 78.3
Table 2: Overall navigation success rate (SR) comparison against baselines. Policies with SRU implicit recurrent memory outperform: (i) GTRL (stacked historical observations) and (ii) EMHP (explicit mapping and historical path). GTRL* denotes our modified GTRL variant where stacked observations are replaced by SRU implicit memory, yielding a substantial gain over the original GTRL baseline.
Figure 12: Average training return rewards for attention ablations (all using SRU recurrent memory): (1) without attention (w/o.) Wijmans et al. (2019); (2) Goal-guided Transformer (GoT) attention Huang et al. (2023); and (3) the proposed two-stage spatial attention (Ours). The proposed spatial attention achieves the highest returns, indicating more effective extraction of task-relevant spatial cues for improved recurrent memorization.
Attention Configuration SR (%)
w/o. Attention 50.5
GoT (GTRL*) 68.4
Spatial Attention (Ours) 78.9
Table 3: Navigation success rate (SR) for policies integrated with different attention configurations. The policy with the proposed spatial attention (Ours) achieves the highest SR, outperforming (i) the baseline without attention Wijmans et al. (2019) and (ii) the Goal-guided Transformer (GoT) Huang et al. (2023) integrated with SRU memory (GTRL* approach in Sec. 5.3), highlighting the importance of an attention mechanism for training mapless navigation end-to-end and the effectiveness of the proposed spatial attention structure.
@@ 5.4 Importance of Spatial Attention Layers
We now examine the role of the proposed spatial attention layers in the network architecture and evaluate their impact on navigation performance. These layers are designed to compress and emphasize relevant features from encoded observations, addressing a key challenge faced by recurrent structures: the difficulty of retaining long-term information due to the exponential decay of memory over time. By selectively focusing on the most salient features, the attention mechanism emphasizes the most relevant spatial cues for navigation based on the robot’s state and reduces the information density passed into the recurrent memory at each step. We hypothesize that this mechanism can improve the network’s memorization and navigation capabilities, enabling it to handle complex, long-range tasks more effectively.
To test this, we conduct an ablation study by: (i) removing the attention layers from our network architecture and replacing them with convolution followed by average pooling for feature compression, as implemented in Wijmans et al. (2019), and (ii) comparing the performance of our proposed spatial attention layers against the Goal-guided Transformer (GoT) architecture proposed in Huang et al. (2023). The GoT architecture utilizes a modified Vision Transformer (ViT) that integrates the goal state as an additional token. It performs self-attention across both visual feature tokens and the goal token to extract goal-relevant features. In contrast, our approach first applies self-attention exclusively to visual tokens to enhance spatial features. Subsequently, the goal and proprioceptive state are used as queries in the cross-attention layer to compress and extract the most relevant features. For a fair comparison in the ablation experiments, we use identical training settings for all approaches, integrating the SRU memories and the pretrained encoder while varying only the attention layers used to process visual features during RL. The GoT integrated with SRU is the the GTRL* approach, as described in Sec. 5.3. Figure 12 shows the average return rewards during training. The network without the attention layers exhibits significantly lower performance compared to the two policies utilizing attention mechanisms. Additionally, our proposed spatial attention layers outperform the GoT attention mechanism. Table 3 shows the SR performance of the three configurations: (i) without attention, (ii) with GoT attention, and (iii) with our proposed spatial attention (Ours). Our method achieves a 56.2% relative SR improvement over the no-attention baseline, highlighting the importance of selectively compressing and extracting spatial features for long-range mapless navigation when utilizing implicit recurrent memory. Furthermore, it achieves a 15.4% relative improvement (18.1% when trained with the ANYmal robot model, as shown in Table 2) over the policy utilizing GoT attention. This demonstrates that the proposed two-stage spatial attention mechanism more effectively extracts task-relevant cues, enhancing recurrent memorization and policy optimization.
@@
Notably, the attention effect emerges naturally during the end-to-end RL training without requiring additional supervision or auxiliary losses. Figure 11 illustrates the attention weights generated by the cross-attention layer over raw visual inputs in three distinct real-world deployment scenarios: an indoor office, an outdoor terrace, and a forest environment. The attention weights, with four attention heads (depicted in different colors), dynamically emphasize the most relevant spatial cues, such as obstacles and navigable free space, based on the robot’s state at the time the depth input was recorded. This highlights the effectiveness of training the spatial attention mechanism end-to-end and its ability to generalize across diverse and challenging environments.
@@ 5.6 Large-Scale Pretraining for Sim-to-Real Transfer
Table 4: Comparison of the overall navigation success rate (SR) with and without DML regularization for policies using LSTM and SRU units. DML significantly enhances SR for SRU (over 20%) and provides a marginal improvement for LSTM (2.8%), showcasing DML’s effectiveness in unlocking SRU’s potential for long-range mapless navigation.
@@
Figure 18: Evaluation of long-range mapless navigation in diverse real-world environments: (a) Main Hall of a university, (b) outdoor terrace, and (c) forest environment with natural obstacles. In each scenario, the robot is tasked with two separate navigation goals (memory reset between goals), resulting in two trajectory segments (orange and blue). Labels A, B, and C mark key areas traversed by the robot.
@@ 6 Limitations and Future Work
While the proposed SRUs in this paper demonstrate significant improvements in spatial-temporal learning capabilities, their recurrent nature remains subject to exponential memory decay, which can limit their ability to retain global context over extended sequences. As a result, the long-range navigation capabilities presented in this paper are centered on local, mapless navigation using egocentric sensing. In this context, ”long-range” refers to planning horizons that extend well beyond the local perception radius (e.g., 10 m), enabling rerouting from local dead ends without reliance on an explicit global map. Extending this approach to global-scale navigation—spanning kilometers or hours—would likely require additional mechanisms or architectural enhancements, such as the integration or maintenance of a global map.
Furthermore, while SRUs enhance the network’s capacity for implicit spatial memorization and improve long-range navigation performance, the precise characteristics of the information retained and utilized during end-to-end navigation training remain unclear. This highlights a broader challenge in explainable artificial intelligence, where understanding the internal representations and decision-making processes of neural networks continues to be an active area of research (Mi et al., 2024). Future work could explore integrating SRUs with recent advancements in foundation pretraining, such as DINO (Caron et al., 2021), to combine their strengths in scene understanding with the efficiency of recurrent structures, further enhancing the policy’s performance in complex real-world environments. Investigating auxiliary losses or additional regularization techniques to further leverage the potential of spatial-temporal memorization in SRUs could also be beneficial. Additionally, extending the application of SRUs to other domains, such as robotic manipulation and 3D reconstruction, could unlock new possibilities and advancements in spatial-temporal learning. In summary, while SRUs are effective, they represent a simple yet practical solution—not necessarily unique or optimal—that proves successful in our end-to-end mapless navigation context. More importantly, this work aims to highlight the potential of implicit spatial memory mechanisms in addressing complex navigation challenges while identifying opportunities for further exploration and optimization in both methodology and application domains.